OpenStudy (anonymous):

Write and equation of the line that passes through the given point P and has the given slope. P(0,5) slope=2

OpenStudy (anonymous):

I would start by writing the Point-Slope form of a line. \[y-y1=m(x-x1)\]

OpenStudy (anonymous):

Then we can plug in the known values which are slope=m=2, and then a point (x1,y1)=(0,5).

OpenStudy (anonymous):

So we should get... y-5=2(x-0) ...simplifying y-5=2(x) y-5=2x y=2x+5

OpenStudy (anonymous):

Hope that helps :)

OpenStudy (anonymous):

Thank you very helpful.

OpenStudy (anonymous):

Thank you.